OpenStudy (kainui):

Medal+fan to whomever finds the most solutions/best ideas about this problem for fun:

OpenStudy (kainui):

\[\LARGE \frac{d^n}{dx^n}(y^m)=\frac{d^m}{dx^m}(y^n)\] what are y, n, and m's we can choose to get a true statement here?

OpenStudy (freckles):

one obvious one is m=n=1

OpenStudy (freckles):

or just m=n

OpenStudy (anonymous):

http://www.silvergames.com/trollface try this @Kainui

OpenStudy (confluxepic):

I remember solving a similar question.

OpenStudy (anonymous):

lol

OpenStudy (freckles):

y=e^x for any m or n is one solution m=n for any y is another solution thinking on more

OpenStudy (anonymous):

dang @freckles are you like a genius or something lol

OpenStudy (freckles):

y=c for any m or n where c is a constant

OpenStudy (freckles):

No I'm not a genius

OpenStudy (anonymous):

yes u are lol

OpenStudy (anonymous):

yeah i agree with @magepker728 cause thats like chemistry lol idk or that hard science crap lol, and you didnt get confused on it so yeah thats proof

OpenStudy (anonymous):

how do you not have a title like @magepker728

OpenStudy (freckles):

and i do have to make a correct I didn't notice the y where raised

OpenStudy (freckles):

y=e^x where m=n is a solution*

OpenStudy (anonymous):

pshh still lol

OpenStudy (kainui):

Here's one: \[\LARGE \frac{d^2}{dx^2}(\tan x)=\frac{d}{dx}(\tan^2 x) \] where m=1, n=2, and y=tan(x) =P

OpenStudy (freckles):

\[\frac{d^2}{dx^2}(-\cot(x))=\frac{d}{dx}\csc^2(x)=-2\csc^2(x)\cot(x) \\ \frac{d}{dx}\cot^2(x)=-2\cot(x)\csc^2(x)\]

OpenStudy (freckles):

so where m=1,n=2, and y=-cot(x)

OpenStudy (freckles):

Kinda stole that one from you though